OpenStudy (dls):

[CHALLENGE]The differential co-efficient of f(log x),where f(x)=logx is?[CHALLENGE] A)x/logx B)(xlogx)^-1 C)logx/x D)None of these

OpenStudy (dls):

@hartnn

OpenStudy (dls):

its not easy btw..

OpenStudy (dls):

@sirm3d

OpenStudy (dls):

@yrelhan4

OpenStudy (yrelhan4):

df(x)/dx nikalna hai?

OpenStudy (dls):

no idea

OpenStudy (dls):

im very confused and thisis a tricky question,take it as an exception

OpenStudy (yrelhan4):

pehle vo dekh kya hota hai. aur fir bta. kyunki mereko bhi nhi pta. :P but shayad df(x)/dx hi hota hai.

OpenStudy (dls):

asksomeone :|

OpenStudy (yrelhan4):

huh :/ mai jaa rha hun. tu dekh liyo.

hartnn (hartnn):

its just the derivative.

OpenStudy (dls):

that is what you think,im pretty sure you'll give a wrong answer :P

hartnn (hartnn):

then i have nothing to do here :P

OpenStudy (dls):

lol atleast make a guess

OpenStudy (dls):

A/B/C/D

hartnn (hartnn):

derivative : f(log x) = log (log x) f'(log x) = (1/ log x ) (log x)' = 1/ (x log x) = B

OpenStudy (dls):

haha :| wrong

hartnn (hartnn):

:'P

OpenStudy (anonymous):

I think your math teacher is crazy

OpenStudy (anonymous):

I looked up the definition of differential coefficient on other sites as well . . . they all said the same thing

OpenStudy (kropot72):

The answer is B)

OpenStudy (kropot72):

If y = log f(x) \[\frac{dy}{dx}=\frac{f \prime (x)}{f(x)}\]

OpenStudy (kropot72):

f(x) = log x f(log x) = log (log x) y = log (log x) f(x) = log x f'(x) = 1/x \[\frac{dy}{dx}=\frac{f \prime (x)}{f(x)}=\frac{1}{x \log_{} x}\]

OpenStudy (dls):

its D :)

OpenStudy (anonymous):

i can answer it

OpenStudy (dls):

go ahead ! :)

OpenStudy (anonymous):

I still think you teacher is crazy . . . what does he teach you differential co-efficient means?

OpenStudy (anonymous):